A local pharmacy manager decides to raise the price of a 50-pill prescription of

amoxicillin (an antibiotic) from Rs. 1000 to Rs. 1200. The pharmacy tracks the sales

of amoxicillin over the next month and finds that sales decline from 2,000 to 1,980

boxes. Is the demand elastic or Inelastic?


1
Expert's answer
2021-05-03T10:55:42-0400

The price elasticity of demand is:

"Ed = \frac{1,980 - 2,000} {1,200 - 1,000} ×\frac{1,200 + 1,000} {1,980 + 2,000} = - 0.055."

So, the demand is inelastic.